# Towards Full-scene Domain Generalization in Multi-agent Collaborative Bird’s Eye View Segmentation for Connected and Autonomous Driving

### v2

@@ V Conclusion
In this paper, we have developed a novel framework to tackle the domain generalization with collaborative perception systems for connected and autonomous driving. We first propose an amplitude augmentation method to augment the original source domain data to different target domains in collaboration with our constructed dataset. Then, we have designed a meta-consistency training scheme based on meta-learning paradigm. This training scheme can simulate the domain shift, guiding the model to learn how to learn from different domains, and the meta-consistency loss can force the model to learn the domain-invariant features, thereby enhancing the ability of generalization of the model. Finally, we have leveraged the intra-system domain alignment method to minimize the domain gap among the data perceived by different collaborative vehicles prior to inference. Comprehensive experiments have demonstrated the superiority of our framework compared with the existing state-of-the-art methods.

### v3

@@ VI Conclusion
In this paper, we have developed a novel framework to tackle the domain generalization with collaborative perception systems for connected and autonomous driving. To achieve this goal, we first propose an amplitude augmentation method to transform data from original source domain to various target domains in collaboration with our constructed dataset. Then, we design a meta-consistency training scheme based on meta-learning paradigm, which can simulate domain shifts, guiding the model to learn how to learn from different domains, and the meta-consistency loss can force the model to learn the domain-invariant features, thereby enhancing the ability to generalize. Finally, we leverage the intra-system domain alignment to minimize the domain gap among the data perceived by different CAVs prior to inference. Comprehensive experiments have demonstrated the superiority of our framework compared with the existing state-of-the-art methods.